Solve.
A garden area is 30 feet long and 20 feet wide. A path of uniform width is set around the edge. If the remaining garden area is 400 feet squared, what is the width of the path?

Answer by jaymore(5)

Let the width of the path be x
New dimension will be:
30-2x and 20-2x
Area is 400
(30-2x)(20-2x)=400
Expand to get 4x2-100x+200=0
Divide by 4 throught the equation to get
x2-25x+50=0
Factorize or use quadratic formula get the value of X.